### v1

# Riemannian Time Warping: Multiple Sequence Alignment in Curved Spaces

Abstract
Temporal alignment of multiple signals through time warping is crucial in many fields, such as classification within speech recognition or robot motion learning. Almost all related works are limited to data in Euclidean space. Although an attempt was made in 2011 to adapt this concept to unit quaternions, a general extension to Riemannian manifolds remains absent. Given its importance for numerous applications in robotics and beyond, we introduce Riemannian Time Warping (RTW). This novel approach efficiently aligns multiple signals by considering the geometric structure of the Riemannian manifold in which the data is embedded. Extensive experiments on synthetic and real-world data, including tests with an LBR iiwa robot, demonstrate that RTW consistently outperforms state-of-the-art baselines in both averaging and classification tasks.
Index Terms:
AI-Based Methods, Learning from Demonstration, Multiple Sequence Alignment, Riemannian Manifold
I Introduction
Time warping describes the temporal alignment of several signals – a fundamental operation for processing sequential data commonly used in speech recognition [1, 2], human action recognition [3, 4], or robot motion learning [5, 6]. Given multiple signals with varying lengths describing the same underlying process, each signal is warped with the goal to improve their temporal alignment. Identifying these warping parameters belongs to the domain of unsupervised learning.
Dynamic Time Warping (DTW) [7], introduced in 1978, finds the optimal alignment analytically through Dynamic Programming [8], and is a prominent approach when dealing with just two signals, however, its computational complexity scales exponentially, which prevents its application to larger datasets. For this reason, alternative algorithms have been proposed that sacrifice global optimality to accelerate the search process by several orders of magnitude [9, 10, 11].
Almost all related works focus on data in the Euclidean space. However, many real-world problems describe temporal phenomena evolving on Riemannian manifolds (cf. Figure LABEL:fig:overviewRTW). This is especially important in robotics, where many fundamental problems involve non-Euclidean geometry [12]. Typical examples in robotics are the unit sphere $\mathbb{S}^{3}$ for representing orientations or symmetric positive definite matrices $\mathbb{S}^{D}_{++}$ for representing stiffness gains or manipulability ellipsoids. Taking into account the data structure is important in data-driven machine learning and can be seen as prior geometric knowledge. In this line, Quaternion Dynamic Time Warping (QDTW) [13] extends its predecessor DTW to signals of unit quaternions, but suffers from the same computational complexity. To the best of our knowledge, so far no generic time warping method has been reported that can handle structured non-Euclidean data in general. One reason for this could be that the sum and scalar multiplication operations used in related works are not defined on Riemannian manifolds [14].
As our main contribution, we present a novel approach called Riemannian Time Warping (RTW) for efficient temporal alignment of multiple signals with varying lengths. Our method, for the first time, explicitly considers the Riemannian manifold ${\mathcal{M}}$ on which the data resides. This letter presents an extensive evaluation on challenging benchmarks, involving datasets in $\mathbb{R}^{1}$, $\mathbb{R}^{2}$, $\mathbb{S}^{1}$, $\mathbb{S}^{3}$, and $\mathbb{S}^{2}_{++}$, as well as problems with multiple signals capturing complex patterns. In averaging and classification tasks, RTW consistently outperforms the state-of-the-art [10, 11] w.r.t. alignment quality and mean signal representation. Furthermore, RTW achieves near-optimal results compared to techniques based on Dynamic Programming such as [13], however, with significantly reduced computation time, as its computational complexity scales linearly with the number of signals and data points. We also show that learning movement primitives benefits from data-preprocessing with RTW. Finally, we validate our approach in the context of real-robot motion learning with a collaborative torque-controlled KUKA LBR iiwa manipulator, leveraging the Riemannian manifold associated with rigid body motions in $\mathrm{SE}(3)$. For conducting all of these comparisons, we also extended state-of-the-art methods. An accompanying video of this work is available at https://youtu.be/v_k2LwEDH9I.
Notation: In the remainder, we will use square bracket notation to describe discrete values, e.g, ${\mathbf{x}}[t]\in\mathbb{R}^{D},{\bm{x}}[t]\in{\mathcal{M}}$, and round bracket notation for continuous functions, e.g, ${\mathbf{\phi}}(z)$.
II Related Work
Many related works focus on aligning ${N}=2$ signals of length ${T}$. Despite its age, the DTW approach [7] remains relevant and is widely used due to its simplicity and effectiveness. Extending DTW to ${N}>2$ signals introduces significant computational challenges, i.e. refer to Multiple Multi-Dimensional Dynamic Time Warping (MMDDTW) [15], which computes the distance matrix over all ${N}$ dimensions. While this results in a globally optimal solution through Dynamic Programming, it becomes unfeasible for $N\gg 2$ due to its computational complexity of $\mathcal{O}({T}^{{N}})$. The Non-Linear Alignment and Averaging Framework (NLAAF) [16] is another extension that iteratively aligns pairs of signals while adjusting previous solutions, with a complexity of $\mathcal{O}({N}{T}^{2})$.
Generalized Time Warping (GTW) [9], Trainable Time Warping (TTW) [10], and Neural Time Warping (NTW) [11] are newer methods that align many signals in linear time. GTW uses a pre-defined set of basis functions, which is cumbersome for complex warpings as many basis functions must be defined beforehand. TTW and NTW, on the other hand, first transform the discrete signals into the continuous time domain through sinc interpolation, apply an index shift in the continuous space, and finally sample the shifted signal back into the discrete time domain. TTW models the warpings by estimating coefficients of the discrete sine transform, while NTW uses a neural network. Both algorithms outperform GTW [10, 11], therefore we do not consider GTW in our experiments.
To the best of our knowledge, QDTW [13] is the only warping scheme that explicitly deals with a Riemannian manifold, focusing exclusively on unit quaternions. However, in recent years, Riemannian manifolds have gained popularity in various domains – especially in robotics [17, 18, 19, 20], as they provide a convenient way to generalize methods originally developed for the Euclidean space. This letter presents the first approach for time warping on generic Riemannian manifolds, that applies to datasets with ${N}\gg 2$. Table I compares related works.
TABLE I: Comparison between state-of-the-art Time Warping Methods
[…]
DTW [7] = $=\,2$$\times$ $\times$ $\mathcal{O}({T}^{2})$
QTDW [13] = $=\,2$$\times$ $\mathcal{O}({T}^{2})$
MMDDTW [15] > $>\,2$$\times$ $\times$ $\mathcal{O}({T}^{N})$
TTW [10] $\gg 2$$\times$ $\times$ $\mathcal{O}({N}{Z})$
NTW [11] $\gg 2$$\times$ $\times$ $\mathcal{O}({N}{Z}{T})$
RTW [proposed] $\gg 2$ $\mathcal{O}({N}{Z})$
III Time Warping in the Continuous Time Domain
Here we first introduce a simplified version of our approach dedicated to Euclidean data ${\mathbf{x}}\in\mathbb{R}^{D}$. The full Riemannian Time Warping (RTW) algorithm considering curved spaces with ${\bm{x}}\in{\mathcal{M}}$ is then proposed in the next section.
III-A Problem Statement
Consider ${N}$ discrete signals ${\mathbf{X}}=\{{\mathbf{x}_{n}}\}_{n=1}^{N}$ of lengths ${T_{n}}$, i.e. ${\mathbf{x}_{n}}=\{{\mathbf{x}_{n}}[t]\}_{t=1}^{{T_{n}}}$, where ${\mathbf{x}_{n}}[t]\in\mathbb{R}^{D}$. Signals with temporal modifications ${\mathbf{{\hat{X}}}}=\{{\mathbf{{\hat{x}}}_{n}}\}_{n=1}^{N}$ of new length ${Z}\geq{T_{\text{max}}}=\max{\left\{{T_{n}}\right\}}_{n=1}^{{N}}$, with ${\mathbf{{\hat{x}}}_{n}}=\{{\mathbf{{\hat{x}}}_{n}}[z]\}_{z=1}^{Z}$ and ${\mathbf{{\hat{x}}}_{n}}[z]\in\mathbb{R}^{D}$ are obtained by applying a set of parameterized warping functions ${\mathbf{\Omega}}={\left\{{\mathbf{\omega}_{n}}[z]:{\left\{1,\cdots,Z\right\}}\mapsto[0,1]\right\}}_{n=1}^{N}$. The Multiple Sequence Alignment (MSA) problem consists of
[…]
where the warping functions ${\mathbf{\omega}_{n}}$ must satisfy three constraints
Boundary Constraint: ${\mathbf{\omega}_{n}}[1]=0,{\mathbf{\omega}_{n}}[Z]=1$ (2a)
Continuity Constraint: ${\mathbf{\omega}_{n}}[z+1]-{\mathbf{\omega}_{n}}[z]\leq\frac{1}{{T_{\text{max}}}}$ (2b)
Monotonicity Constraint: ${\mathbf{\omega}_{n}}[z]\leq{\mathbf{\omega}_{n}}[z+1]$ (2c)
and ${\mathcal{D}_{{\mathbf{\Omega}}}({\mathbf{{\hat{X}}}})}$ is a distance measure between all warped signals ${\mathbf{{\hat{x}}}_{n}}$ and their mean
${\mathbf{{\hat{y}}}}=\frac{1}{N}\textstyle\sum_{n=1}^{N}{\mathbf{{\hat{x}}}_{n}}\,.$ (3)
The challenge is to find optimal and feasible warping functions efficiently. This is non-trivial, as ${\mathbf{X}}$ is a set of discrete signals.
III-B Riemannian Time Warping (RTW) Algorithm in $\mathbb{R}^{D}$
We transform ${\mathbf{X}}$ into the continuous time domain to compute the aligned signals ${\mathbf{{\hat{X}}}}$ through sinc interpolation
${\mathbf{{\hat{x}}}_{n}}[z]=\sum_{m=\lfloor{{\mathbf{\omega}_{n}}[z]{T_{n}}}\rfloor{}-{\nu}}^{\lfloor{{\mathbf{\omega}_{n}}[z]{T_{n}}}\rfloor{}+{\nu}}{\mathbf{x}_{n}}[m]\operatorname{sinc}(m-{\mathbf{\omega}_{n}}[z]\,\,{T_{n}})\,,$ (4)
[…]
${\mathbf{\omega}}_{n}[z]=\hat{z}\sqrt{{N}}{\mathbf{e}_{1}}+\hat{z}{\left(1-\hat{z}\right)}\sum_{k=1}^{N-1}{\left[{\mathbf{\phi}_{\theta}}\left(\hat{z}\right)\right]}_{k}\,{\mathbf{e}_{k+1}}\,,$ (5)
where the $\theta$-parameterized function ${\mathbf{\phi}_{\theta}}(\hat{z}):{\left[0,1\right]}\mapsto\mathbb{R}^{{N}-1}$ is modeled as a neural network, and $\hat{z}=\frac{z-1}{Z-1}$. The boundary constraint (2a) is guaranteed by design. During warping, indices of a signal may be duplicated to improve the temporal alignment, requiring a longer signal, i.e, $Z\geq{T_{\text{max}}}\,$. Hence, the continuity constraint (2b) is fullfilled with $Z={N}{T_{\text{max}}}$ [11]. To enforce the monotonicity constraint (2c), a ${\lambda}$-weighted penalty term is added to the loss as in [11]
${\mathbf{\Omega}}_{opt}=\operatorname*{argmin}_{{\mathbf{\Omega}}}{\left({\mathcal{D}_{{\mathbf{\Omega}}}({\mathbf{{\hat{X}}}})}+{\lambda}\,{p({\mathbf{\Omega}})}\right)}\,,$ (6)
${p({\mathbf{\Omega}})}=\sum_{n=1}^{{N}}\sum_{z=1}^{Z-1}\max{\left({\mathbf{\omega}_{n}}[z]-{\mathbf{\omega}_{n}}[z+1]\,,0\right)}$ (7)
The unconstrained optimization problem (6) approximates (1) and is solved using state-of-the-art optimization techniques.
Finally, we define the distance measure ${\mathcal{D}_{{\mathbf{\Omega}}}({\mathbf{{\hat{X}}}})}$ as a Gaussian-weighted window loss of point segments
$\begin{gathered}{\mathcal{D}_{{\mathbf{\Omega}}}({\mathbf{{\hat{X}}}})}=\frac{1}{{N}{Z}}\sum_{n=1}^{{N}}\sum_{z=1}^{{Z}}\sum_{k=-{\kappa}}^{{\kappa}}{\hat{g}(v,z)}\,{\text{dist}({\mathbf{{\hat{x}}}_{n}}[v],{\mathbf{{\hat{y}}}}[v])}\,,\\ {\hat{g}(v,z)}=\frac{{g(v)}}{\sum_{i=-{\kappa}}^{{\kappa}}{g(z+i{\rho})}}\,\,\,\,\,,\,\,\,\,\,{g(j)}=\exp{\left(-\frac{j^{2}}{2\sigma^{2}}\right)}\,,\end{gathered}$ (8)
with $v=z+k{\rho}$, where the window size ${\kappa}\in\mathbb{N}_{0}$ describes how many additional indices are considered at index $z$ and the step size ${\rho}\in\mathbb{N}$ defines the distance between them, resulting in the indices $v$ that form a segment [^4] As for the sinc interpolation, we apply clamping to always receive valid indices for a full segment, i.e., we set $v=\max(1,\min({Z},v))$. . Each index within a segment is weighted with the Gaussian kernel ${\hat{g}(v,z)}$. Given the empirical rule, we set $\sigma=\frac{{\kappa}{\rho}}{3}+\epsilon$ with a small $\epsilon>0$ for numerical reasons to ensure a weighting, such that all indices lie within $99,7\%$ of the Gaussian’s power, emphasizing central indices as visualized in Figure -1. Note that by setting ${\kappa}=0$, we receive a point-to-point metric similar to TTW or NTW. For data in $\mathbb{R}^{D}$, we utilize the Euclidean distance metric
$\text{dist}_{\mathbb{R}^{D}}({\mathbf{{\hat{x}}}}[z],{\mathbf{{\hat{y}}}}[z])={\left\|{\mathbf{{\hat{x}}}}[z]-{\mathbf{{\hat{y}}}}[z]\right\|}\,.$ (9)
Figure -1: Visualization of the Gaussian-weighted window loss between two signals ${\mathbf{{\hat{x}}}}$ (green) and ${\mathbf{{\hat{y}}}}$ (red) as defined in (8). The example shows one segment at index $z=50$ with window size ${\kappa}=4$ and step size ${\rho}=5$, where red, yellow and green mark the areas covered by one, two and three times the variance $\sigma$ of the Gaussian. By considering a segment around index $z$, the first peak of both signals becomes distinguishable from the second.
IV RTW Extension for Riemannian Manifolds
Extending RTW to data on a Riemannian manifold ${\mathcal{M}}$ is not straightforward as there is no sum and scalar multiplication defined, i.e., it is not a vector space. Accordingly, the classical time warping approaches yield impractical and nonsensical results, because the warped signals and their mean would not reside on the manifold. Specifically, sinc interpolation (4), mean computation (3) and distance evaluation (9) would provide outcomes that do not respect the manifold structure. To address this, our key idea is to utilize Euclidean tangent spaces ${\mathcal{T}}$ that locally preserve distances. Then, we adjust the sinc interpolation (see IV-A), mean computation (see IV-B) and distance evaluation (see IV-C), to incorporate the curved manifold structure by operating within the tangent spaces. These novel modifications enable temporal alignment in the continuous time domain to be applied to Riemannian manifolds without affecting the computational complexity $\mathcal{O}({N}{Z})$: RTW scales linearly with ${N}$ and ${Z}$.
More specifically, a manifold ${\mathcal{M}}$ is a $D$-dimensional smooth space [14], for which there exists a tangent space ${\mathcal{T}_{{\bm{e}}}}$ for each point ${\bm{e}}\in{\mathcal{M}}$ on the manifold, with ${\bm{e}}$ being the origin of ${\mathcal{T}_{{\bm{e}}}}$. A manifold with a positive definite inner product defined on the tangent space is called Riemannian. The exponential map ${\mathbf{Exp}_{{\bm{e}}}({\bm{u}})}:{\mathcal{T}_{{\bm{e}}}}\mapsto{\mathcal{M}}$ maps a point ${\bm{u}}\in{\mathcal{T}_{{\bm{e}}}}$ from the tangent space onto the manifold. The inverse operation is the logarithmic map ${\mathbf{Log}_{{\bm{e}}}({\bm{x}})}:{\mathcal{M}}\mapsto{\mathcal{T}_{{\bm{e}}}}$. These mappings are distance-preserving, i.e. the geodesic distance between ${\bm{x}}$ and ${\bm{e}}$ is identical to ${\left\|{\mathbf{Log}_{{\bm{e}}}({\bm{x}})}\right\|}$. This feature allows indirect computation to be performed on the manifold by applying Euclidean methods in the tangent space. The formulation of the exponential and logarithmic map is manifold dependent.
IV-A Windowed Sinc Interpolation on Riemannian Manifolds
The windowed sinc interpolation in (4) for a time index $z$ constitutes a weighted sum, which is not a closed operation under ${\mathcal{M}}$. Therefore, we project the original $12{\nu}+$ 1 data points from the manifold onto the tangent space ${\mathcal{T}_{{\bm{e}_{n}}[z]}}$ through ${\bm{u}_{n,z}}[m]={\mathbf{Log}_{{\bm{e}_{n}}[z]}({\bm{x}_{n}}[m])}$, where we warp and interpolate
${\bm{{\hat{u}}}_{n}}[z]=\sum_{m=\lfloor{{\mathbf{\omega}_{n}}[z]{T_{n}}}\rfloor{}-{\nu}}^{\lfloor{{\mathbf{\omega}_{n}}[z]{T_{n}}}\rfloor{}+{\nu}}{\bm{u}_{n,z}}[m]\operatorname{sinc}(m-{\mathbf{\omega}_{n}}[z]\,{T_{n}})\,.$ (10)
Then, the result for the time index $z$ is projected from the specific tangent space ${\mathcal{T}_{{\bm{e}_{n}}[z]}}$ back onto the manifold via
${\bm{{\hat{x}}}_{n}}[z]={\mathbf{Exp}_{{\bm{e}_{n}}[z]}({\bm{{\hat{u}}}_{n}}[z])}\,.$ (11)
Note that the choice of the tangent point ${\bm{e}_{n}}[z]\in{\mathcal{M}}$ is important to minimize undesired distortions [12]. Defining the origin of the tangent space at the center of the interpolation window as ${\bm{e}_{n}}[z]={\bm{x}_{n}}{\left[\,\lfloor{{\mathbf{\omega}_{n}}[z]\,{T_{n}}}\rfloor{}\,\right]}$, implies least distortion for points ${\bm{x}_{n}}[m]$ with the highest power of the $\operatorname{sinc}$ during interpolation. Further note that by considering a small window size $12{\nu}+$ 1 the distortion effect becomes neglectable. Obviously, the particular tangent space is reevaluated for each time index $z$ and signal $n$.
IV-B Signal Mean Computation on Riemannian Manifolds
We need to adjust the mean computation of the warped signals (3), as it is not a closed operation under ${\mathcal{M}}$. Starting from an initial estimate ${\bm{{\hat{y}}}}[z]\in{\mathcal{M}}$, we utilize Gauss-Newton iterations by projecting the $z$-th data point ${\left\{{\bm{{\hat{x}}}_{n}}[z]\in{\mathcal{M}}\right\}}_{n=1}^{{N}}$ of each signal onto the tangent space ${\mathcal{T}_{{\bm{{\hat{y}}}_{n}}}}$ through ${\bm{{\hat{u}}}_{n}}[z]={\mathbf{Log}_{{\bm{{\hat{y}}}}[z]}({\bm{{\hat{x}}}_{n}}[z])}$. We obtain the center of these projections as
${\bm{{\hat{v}}}}[z]=\frac{1}{{N}}\sum_{n=1}^{{N}}{\bm{{\hat{u}}}_{n}}[z]\,.$ (12)
This center is projected back from the tangent space ${\mathcal{T}_{{\bm{{\hat{y}}}}[z]}}$ onto the manifold ${\mathcal{M}}$ to update the mean at time index $z$ through
${\bm{{\hat{y}}}}[z]\leftarrow{\mathbf{Exp}_{{\bm{{\hat{y}}}}[z]}({\bm{{\hat{v}}}}[z])}\,.$ (13)
This process, (12) and (13), is repeated until convergence, which is typically reached after few iterations, e.g., see [12].
IV-C Distance Computation on Riemannian Manifolds
The distance metric (9) is adapted as well. We utilize the geodesic distance that applies to all Riemannian manifolds
$\text{dist}_{{\mathcal{M}}}({\mathbf{{\hat{x}}}}[z],{\mathbf{{\hat{y}}}}[z])={\left\|{\mathbf{Log}_{{\bm{{\hat{y}}}}[z]}({\bm{{\hat{x}}}}[z])}\right\|}\,.$ (14)
IV-D Further Remarks on the Overview provided in Fig. LABEL:fig:overviewRTW
Figure LABEL:fig:overviewRTW provides a schematic overview, summarizing the proposed RTW approach. Three signals with significant temporal shifts are considered in $\mathbb{S}^{2}$, a unit sphere in $\mathbb{R}^{3}$. The color gradient from blue to red indicates the progression of time along the trajectories. We illustrate the mapping of one signal onto the tangent space for two different origins ${\bm{e}_{n}}[z]$, where the tangent space is a two-dimensional plane. As expected, distortion is noticeable only for data points far from the origin. For visualization purposes, we choose a large interpolation window ${\nu}$, indicated by the bold segments. Note that these parts are close to the origin of the tangent space but not spatially centered around it, as the window is centered in the time domain instead. Also, RTW only projects data points from a single signal onto each tangent space during interpolation, while related works often project all data points onto them, e.g., see [21, 12, 19]. Applying the warping functions ${\mathbf{\omega}_{n}}$ does not alter the shape of a signal in the tangent spaces; rather, the signal is temporally adjusted, as shown by the new color coding. RTW results in three signals on the manifold that are well aligned. After a few Gauss-Newton iterations, we obtain the signal mean, respecting the unit sphere structure.
V Simulations and Experiments
This section compares the proposed RTW approach with baselines in various settings considering datasets in $\mathbb{R}^{1}$, $\mathbb{S}^{1}$, $\mathbb{S}^{3}$ and $\mathbb{S}^{2}_{++}$. We also apply RTW as a pre-processing step for learning movement primitives, on both hand-written data in $\mathbb{R}^{2}$ and within a real-robot scenario in $\mathrm{SE}(3)$ using an LBR iiwa.
RTW, TTW and NTW have been implemented in PyTorch, utilizing the Adam optimizer with a learning rate of $0.01$ and Autograd for gradient computations. The best model over 256 epochs was selected during the comparisons. For RTW, we use a neural network to model ${\mathbf{\phi}_{\theta}}(\hat{z})$ with four fully connected linear layers, ReLU activations and skip connections of size 1-512-512-1025-(N-1), as suggested in [11]. However, unlike NTW, we do not initialize the weights with zero, as preliminary trials showed that this significantly degrades the performance. Instead, we apply symmetry breaking [22], initializing weights with the Xavier initialization method [23].
We use a window size of ${\nu}=10$ for the sinc interpolation in RTW/TTW (10). For the Gaussian-weighted window loss in RTW (8), we choose window size ${\kappa}=5$ and step size ${\rho}=5$, which had neglectable impacts on the runtime. The penalty term (6) for RTW/NTW is weighted with ${\lambda}=100$.
The number of learnable parameters $\theta$ in NTW and RTW is $263680+1025({N}-1)+({N}-1)$, i.e., the number of weights and biases from all layers of the neural network, whereas it is only ${N}{K}$ for TTW due to the warping function
${\mathbf{\omega}}_{n}^{ttw}[z]=\hat{z}+\sum_{k=1}^{{K}}\alpha_{n,k}\sin\left(\pi k\hat{z}\right)\quad,\quad\hat{z}=\frac{z-1}{Z-1}\,,$ (15)
where ${K}\in\mathbb{N}_{+}$ defines how many components of the discrete sine tranform are used to model the warping functions.
We also recommend watching the accompanying video.
V-A Extensive Benchmark on the UCR Time Series Archive
We evaluate the performance of the simplified RTW (as presented in III-B) w.r.t. TTW/NTW in $\mathbb{R}^{1}$. Therefore, we perform DTW averaging and classification tasks on the UCR Time Series Archive [24], replicating the experiments in [10, 11] to enable consistent benchmarking and fair comparisons. This archive consists of $128$ datasets covering various applications; each is pre-split into train ${\mathbf{x}^{\text{train}}_{n,c}}$ and test data ${\mathbf{x}^{\text{test}}_{n,c}}$, containing ${N}\in{\left\{16,\ldots,16800\right\}}$ signals classified into $c\in{\left\{2,\ldots,60\right\}}$ classes, with ${T}\in{\left\{15,\ldots,2844\right\}}$ data points, where most signals are of the same length, i.e., ${T_{1}}=\ldots={T_{{N}}}$.
V-A1 DTW Averaging Task
For all datasets, we perform temporal alignment on the training data ${\mathbf{x}^{\text{train}}_{n,c}}$ for each class $c$ separately, to estimate its mean signal ${\mathbf{{\hat{y}}}^{\text{train}}_{c}}$. We then compute the accumulated DTW distance $\sum_{n=1}^{N}{\mathcal{D}_{dtw}\left({\mathbf{{\hat{x}}}^{\text{train}}_{n,c}},{\mathbf{{\hat{y}}}^{\text{train}}_{c}}\right)}$ [7] for all three methods and perform a paired t-test ($\alpha=0.05$), to find statistically significant differences. Table II-1. lists the results, where each entry describes for how many datasets a method significantly outperformed the other: RTW and NTW are comparable, whereas RTW clearly outperforms TTW.
TABLE II: Statistical Comparison on the UCR Time Series Archive
 1. Averaging 2. Classification
 TTW significantly better than RTW 7.34% 5.47%
 RTW significantly better than TTW 39.84% 54.69%
 No significant difference 52.82% 39.84%
 NTW significantly better than RTW 7.81% 6.25%
 RTW significantly better than NTW 8.59% 35.16%
 No significant difference 83.60% 58.59%
V-A2 Classification Task
We conduct classification tasks on the remaining test data ${\mathbf{x}^{\text{test}}_{n,c}}$. Given the estimated means ${\mathbf{{\hat{y}}}^{\text{train}}_{c}}$ for each class $c$ of a dataset, we predict the class label $\hat{c}_{n}$ for all test signals ${\mathbf{x}^{\text{test}}_{n,c}}$ utilizing nearest centroid classification as $\hat{c}_{n}=\operatorname*{argmin}_{c}\mathcal{D}_{dtw}({\mathbf{x}^{\text{test}}_{n,c}},{\mathbf{{\hat{y}}}^{\text{train}}_{c}})$. The results of the paired t-tests ($\alpha=0.05$) between TTW, NTW and RTW are shown in Table II-2. RTW learns complex non-linear warpings, outperforming both TTW and NTW significantly. Compared to a neural network, modeling the warping functions through the discrete sine transform as in TTW (15) limits their flexibility. Without applying symmetry breaking in NTW, the neural network only learns a linear model, which limits its warping functions to a quadratic structure. Figure V-A2 visualizes the classification performance for all 128 datasets.
Figure 0: Classification performance of TTW (red), NTW (yellow), RTW (green) on all 128 datasets from the UCR Time Series Archive. Each point in the chart represents the TTW/NTW accuracy (x-axis) and RTW accuracy (y-axis). Due to overlapping points, a histogram shows the accuracy-distribution at the top.
V-B Inverted Time Warping in $\mathbb{S}^{1}$
footnotetext: Modified version with an ${N}$-dimensional cost matrix for ${N}>2$. footnotetext: Modified version utilizing (10)-(14) for data in $\mathbb{S}^{1}$. footnotetext: CPU used for the experiments is an AMD® Ryzen 5 3500U. footnotetext: GPU used for the experiments is an NVIDIA GeForce RTX 4080 Super.
Signal alignment is typically performed in an unsupervised setting, making it difficult to evaluate the result w.r.t. the globally optimal solution, as this is usually not available for a dataset with $N\gg 2$ signals [11]. Here, we mitigate this issue through an inverted time warping approach [^9] Warping functions that satisfy constraints (2a)-(2c) can also be used for creating an artificial temporal shift of a given signal. Then, optimal warping functions exist to reverse the tempral shift. , which we apply to data in $\mathbb{S}^{1}$. Exponential and logarithmic map are defined as
${\mathbf{Exp}^{\mathbb{S}^{D}}_{{\bm{e}}}\left({\bm{u}}\right)}={\bm{e}}\cos({\left\|{\bm{u}}\right\|})+\frac{{\bm{u}}}{{\left\|{\bm{u}}\right\|}}\sin({\left\|{\bm{u}}\right\|})\,,$ (16)
${\mathbf{Log}^{\mathbb{S}^{D}}_{{\bm{e}}}\left({\bm{x}}\right)}=\arccos({\bm{e}}^{T}{\bm{x}})\frac{{\bm{x}}-{\bm{e}}^{T}{\bm{x}}{\bm{e}}}{{\left\|{\bm{x}}-{\bm{e}}^{T}{\bm{x}}{\bm{e}}\right\|}}\,.$ (17)
More precisely, starting from a single pre-defined signal ${\mathbf{x}_{\text{original}}}$, we create ${N}$ synthetic trajectories ${\mathbf{X}}=\{{\mathbf{x}_{n}}\}_{n=1}^{N}$ of lengths ${T_{n}}=100$, i.e. ${\mathbf{x}_{n}}=\{{\mathbf{x}_{n}}[t]\}_{t=1}^{{T_{n}}}$, where ${\mathbf{x}_{n}}[t]\in\mathbb{S}^{1}$. This is achieved by generating ${N}$ random warping functions ${\mathbf{\omega}_{n}}$, obtained by uniformly sampling from both (5) and (15) with random parameters, as well as from random spline interpolations of various degrees. Sampled warping functions that do not satisfy constraints (2a)-(2c) are rejected. Then, starting from the original signal ${\mathbf{x}_{\text{original}}}$, the synthetic signals ${\mathbf{X}}$ are obtained through (10)-(11), resulting in a diverse dataset. The goal of this evaluation is to reverse this process, by temporally re-aligning the generated data ${\mathbf{X}}$, i.e., we aim for aligned signals with ${\mathbf{x}_{\text{original}}}\overset{!}{=}{\mathbf{{\hat{x}}}_{n}}\,,\forall\,n\in{\left\{1,\ldots,{N}\right\}}$.
The literature does not offer time warping schemes for ${N}>2$ signals with data on Riemannian manifolds. Hence, we modify and extend QDTW (presented for ${N}=2$ and $\mathbb{S}^{3}$) following the concept of MMDDTW [15]. In addition, we change the implementation of the TTW and NTW algorithms to utilize (10)-(14). These modified algorithms are compared with RTW in two settings, with ${N}=4$ and ${N}=30$ signals. For transparent results, we measure the computation time on a high-grade GPU and a standard CPU. Furthermore, we evaluate three well-established metrics
Restoration Accuracy: $\textstyle\sum_{n=1}^{{N}}{\mathcal{D}_{dtw}\left({\mathbf{{\hat{x}}}_{n}},{\mathbf{x}_{\text{original}}}\right)}$ (18a)
Barycenter Loss: $\textstyle\sum_{n=1}^{{N}}{\mathcal{D}_{dtw}\left({\mathbf{x}_{n}},{\mathbf{{\hat{y}}}}\right)}$ (18b)
Alignment Quality: $\textstyle\sum_{n=1}^{{N}}{\mathcal{D}_{dtw}\left({\mathbf{{\hat{x}}}_{n}},{\mathbf{{\hat{y}}}}\right)}$ (18c)
where the cosine distance $\text{dist}_{\text{cos}}=1-{\left\langle{\mathbf{{\hat{x}}}_{n}}[z],{\mathbf{{\hat{y}}}}[z]\right\rangle}$ is used within each ${\mathcal{D}_{dtw}}$ distance [7]. The Restoration Accuracy describes how effectively random warping functions are reversed to restore the original signal ${\mathbf{x}_{\text{original}}}$. The Barycenter Loss measures how accurately the mean ${\mathbf{{\hat{y}}}}$ of warped signals ${\mathbf{{\hat{x}}}_{n}}$ represents the data ${\mathbf{x}_{n}}$. The Alignment Quality indicates closeness of aligned signals ${\mathbf{{\hat{x}}}_{n}}$ to their mean ${\mathbf{{\hat{y}}}}$. Each experiment is repeated 100 times to compute the average for each metric.
TABLE III: Inverted Time Warping on $\mathbb{S}^{1}$ for ${N}=4$, ${T}=100$
Metric QDTWfootnotemark: TTWfootnotemark: NTWfootnotemark: RTW
Restoration Accuracy 0.6565 0.6785 0.4355 0.2909
Barycenter Loss 0.8399 1.0347 0.6914 0.6932
Alignment Quality 0.0035 0.0341 0.1172 0.0087
CPUfootnotemark: Runtime in [s] 3834.3273 3.3734 10.7157 8.4883
GPUfootnotemark: Runtime in [s] 1505.9386 1.5650 1.6639 1.7574
TABLE IV: Inverted Time Warping on $\mathbb{S}^{1}$ for ${N}=30$, ${T}=100$
Metric QDTWfootnotemark: TTWfootnotemark: NTWfootnotemark: RTW
Restoration Accuracy 2.9523 1.4193 0.3273
Barycenter Loss 6.6766 4.6503 4.4798
Alignment Quality 0.1491 0.8514 0.1095
CPUfootnotemark: Runtime in [s] 18.8754 226.9687 103.1399
GPUfootnotemark: Runtime in [s] 9.4830 10.9298 10.2900
Figure 1: Time warping on the $\mathbb{S}^{1}$ Riemannian manifold. All data points lie on the unit circle (grey dots), however, for visualization purposes we vary the radius to also indicate the time information between 00 and $1$. Starting from an original signal ${\mathbf{x}_{\text{original}}}$ (dashed line), ${N}=30$ random signals (transparent lines) are generated through an inverted time warping approach (see Sec. V-B).
V-B1 Evaluation for ${N}=4$
Table III shows the result. QDTW achieves the best alignment quality due to globally optimal Dynamic Programming. However, RTW is a close second, requiring about three orders of magnitude less computation time. Furthermore, the original signal ${\mathbf{x}_{\text{original}}}$ was restored best by RTW. NTW and RTW perform equally well w.r.t. the Barycenter Loss. The slightly longer runtime of RTW when compared to TTW is justified by its superior results.
V-B2 Evaluation for ${N}=30$
The result is shown in Figure 1 and Table IV. QDTW is not capable of handling such a large number of signals in a reasonable time due to the excessive Dynamic Programming technique. RTW achieves the best result w.r.t. Alignment Quality. Even though TTW also achieves a good alignment, the other two metrics reveal that the signal structure is modified. NTW and RTW are again comparable in terms of the Barycenter Loss, but RTW clearly outperforms both baselines w.r.t. the Restoration Accuracy. The benefits of using the windowed sinc interpolation become evident in the TTW and RTW runtime. Furthermore, the reduced number of parameters optimized within TTW positively affects the CPU runtime. However, on the GPU, all methods are comparably fast for ${T}=100$, due to the advantages of parallel computing.
V-C Multi-dimensional Riemannian manifolds
Next, we extend our experiments to higher-dimensional data by applying RTW to the Riemannian manifolds $\mathbb{S}^{3}$ and $\mathbb{S}^{2}_{++}$.
V-C1 Evaluation in $\mathbb{S}^{3}$
To further validate our results from section V-B, we apply RTW to signals of unit quaternions, as QDTW was originally proposed for data in $\mathbb{S}^{3}$. Similar to [12], we artificially generate ${N}=4$ valid signals of length ${T}=200$ from the LASA [25] dataset. This involves projecting handwritten motion data from $\mathbb{R}^{2}$ onto the unit sphere of $\mathbb{S}^{3}$.
Table V compares QDTW and RTW. We observe that the mean signal obtained from RTW is an improved representation of the original data compared to QDTW. Additionally, RTW achieves an almost optimal alignment quality comparable to QDTW, which is consistent with our previous findings.
TABLE V: Time Warping on $\mathbb{S}^{3}$ for ${N}=4$, ${T}=200$
Metric QDTWfootnotemark: RTW
Barycenter Loss 0.0388 0.0327
Alignment Quality 0.0365 0.0379
V-C2 Evaluation in $\mathbb{S}^{2}_{++}$
We temporally align signals from the Riemannian manifold of symmetric positive definite (SPD) $2\times 2$ matrices. The exponential and logarithmic maps are
${\mathbf{Exp}^{\mathbb{S}^{D}_{++}}_{{\bm{e}}}\left({\bm{u}}\right)}=\sqrt{{\bm{e}}}\exp{\left(\sqrt{{\bm{e}}}^{-1}{\bm{u}}\sqrt{{\bm{e}}}^{-1}\right)}\sqrt{{\bm{e}}},$ (19)
${\mathbf{Log}^{\mathbb{S}^{D}_{++}}_{{\bm{e}}}\left({\bm{x}}\right)}=\sqrt{{\bm{e}}}\log{\left(\sqrt{{\bm{e}}}^{-1}{\bm{x}}\sqrt{{\bm{e}}}^{-1}\right)}\sqrt{{\bm{e}}},$ (20)
where $\sqrt{{\bm{e}}}$ and $\sqrt{{\bm{e}}}^{-1}$ denote the matrix square root and its inverse, respectively. Matrices ${\bm{x}}$ and ${\bm{u}}$ have equal dimensions.
We compare RTW with two modified versions of MMDDTW. The first version, denoted as $\text{MMDDTW}_{{\mathcal{M}}}$, computes the distance between two SPD matrices according to the geodesic distance on ${\mathcal{M}}$ defined in (14) with the Frobenius norm ${\left\|\cdot\right\|}_{F}$. The second version, denoted as $\text{MMDDTW}_{\operatorname{chol}}$, utilizes the Cholesky decomposition $\operatorname{chol}\left(\cdot\right)$ to linearize the manifold, and computes a distance through
$\text{dist}_{\operatorname{chol}}({\mathbf{{\hat{x}}}}[z],{\mathbf{{\hat{y}}}}[z])={\left\|\operatorname{chol}\left({\mathbf{{\hat{x}}}}[z]\right)-\operatorname{chol}\left({\mathbf{{\hat{y}}}}[z]\right)\right\|}_{F}\,.$ (21)
The data for this experiment consists of manipulability ellipsoids, which are artificially generated from the LASA [25] dataset as in [26]. A planar robot with two joints is controlled to follow the handwritten character G as shown in Figure 2. At each time step, the manipulability is computed, resulting in a dataset of ${N}=3$ signals with a length of ${T}=200$ each.
We evaluate the Barycenter Loss (18b) and the Alignment Quality (18c) using both distance measures (14) and (21). Table VI presents the results, supporting our previous findings: RTW outperforms MMDDTW in terms of Barycenter Loss, achieving near-optimal alignment with remarkable speed.
Figure 2: A planar robot (solid lines) follows endeffector trajectories (dashed lines), illustrated at $t=50$ and $t=100$ for each signal respectively (top-left and top-right). The configuration-dependent manipulability is computed for each time step and shown before and after applying RTW (bottom).
TABLE VI: Time warping with SPD matrices for ${N}=3$, ${T}=200$
Metric $\text{MMDDTW}_{{\mathcal{M}}}$ $\text{MMDDTW}_{\operatorname{chol}}$ RTW
$\text{Barycenter Loss}_{{\mathcal{M}}}$ 1.0153 1.0370 0.9579
$\text{Barycenter Loss}_{\operatorname{chol}}$ 0.3449 0.3387 0.3285
$\text{Alignment Quality}_{{\mathcal{M}}}$ 1.0166 1.0463 1.0566
$\text{Alignment Quality}_{\operatorname{chol}}$ 0.3488 0.3389 0.3447
V-D Learning Movement Primitives with RTW pre-processing
Kinesthetic teaching provides an easy-to-use interface to learn robot motion skills. Demonstrated trajectories commonly vary in space and time, and are used to imitate the demonstrated behavior, for example using Dynamic Movement Primitives (DMP) [27], Gaussian Mixture Models (GMM) [6] or Gaussian Processes (GP). As a baseline, the mean signal ${\mathbf{{\hat{y}}}}$ may also serve as time-dependent motion skill encoding. Time warping has been utilized as a pre-processing technique for motion represented in Euclidean space (such as the robot’s joint-space) [5, 6]. However, to the best of our knowledge, related works on end-effector motion in $\mathrm{SE}(3)$ directly learn from raw data without any pre-processing, e.g., [21].
We assess the advantages of using RTW as a pre-processing step rather than relying on raw data, focusing on $\mathbb{R}^{2}$ in this subsection and addressing $\mathrm{SE}(3)$ afterwards. Given ${N}=4$ signals of length ${T}=200$ from the LASA dataset [25], we perform leave one signal-out cross validation.
TABLE VII: Reproduction error of unseen signals with Imitation Learning
DTW distance MAE
Method Raw Data RTW Data Raw Data RTW Data
Mean 0.4693 0.4612 1.4336 1.4292
DMP 0.4861 0.4705 1.4845 1.4560
GMM 0.4820 0.4726 1.4712 1.4682
GP 0.5648 0.5238 1.5949 1.5168
Considering all 26 characters of the dataset results in total 104 experiments. Table VII shows the average DTW distance and mean absolute error (MAE). Across all experiments, applying pre-processing with RTW significantly improves the learning result. Interestingly, the mean signal ${\mathbf{{\hat{y}}}}$ obtained from RTW yields better results than utilizing DMP, GMM and GP in this experiment.
V-E Real-Robot Teaching with Time Warping in $\mathrm{SE}(3)$
Figure 3: Comparison of the results for both robot teaching tasks with naive interpolation, modified TTW/NTW and the proposed RTW approach. The hand-guided demonstration data ${\mathbf{x}_{n}}$ is drawn in grey and the time evolution of the mean signal ${\mathbf{{\hat{y}}}}$ is visualized using the color map. Top: the robot has to follow a constrained path, which requires precise positioning and orientation. The results are shown from a side perspective and the bars on the trajectory indicate the end-effector’s orientation (drawn ten times), which has to be perpendicular to the motion direction. Bottom: the robot has to draw three circles, where the difficulty lies in identifying and aligning the repetitive characteristics correctly. The results are shown from a top-down perspective.
Finally, we apply time warping in real-robot teaching scenarios, utilizing a KUKA LBR iiwa manipulator with seven actuated joints. Taking advantage of the intuitive hand-guiding control mode [28], we record ${N}=4$ signals with varying lengths ${T_{n}}\in\{2048,\ldots,3667\}$ at $30$ Hz. Subsequently, the mean signal ${\mathbf{{\hat{y}}}}$ obtained from temporal alignment is executed employing a state-of-the-art QP-controller [29], taking hardware limits and manipulator dynamics into account.
The trajectory teaching is considered in $\mathrm{SE}(3)$, which is common practice in related works. We denote the end-effector pose ${\bm{x}}=\left[\begin{smallmatrix}{\bm{p}}\\ {\bm{q}}\end{smallmatrix}\right]\in{\left\{\mathbb{R}^{3}\times\mathbb{S}^{3}\right\}}$ consisting of position ${\bm{p}}\in\mathbb{R}^{3}$ and unit quaternion ${\bm{q}}\in\mathbb{S}^{3}$. Similar to [21], we concatenate the logarithmic and exponential map into single operations
${\mathbf{Log}^{\mathrm{SE}(3)}_{{\bm{e}}}\left({\bm{x}}\right)}=\left[\begin{matrix}{\bm{x}_{{\bm{p}}}}\\ {\mathbf{Log}^{\mathbb{S}^{3}}_{{\bm{e}}}\left({\bm{x}_{{\bm{q}}}}\right)}\end{matrix}\right]=\left[\begin{matrix}{\bm{u}_{{\bm{p}}}}\\ {\bm{u}_{{\bm{q}}}}\end{matrix}\right]={\bm{u}}\,,$ (22)
${\mathbf{Exp}^{\mathrm{SE}(3)}_{{\bm{e}}}\left({\bm{u}}\right)}=\left[\begin{matrix}{\bm{u}_{{\bm{p}}}}\\ {\mathbf{Exp}^{\mathbb{S}^{3}}_{{\bm{e}}}\left({\bm{u}_{{\bm{q}}}}\right)}\end{matrix}\right]=\left[\begin{matrix}{\bm{x}_{{\bm{p}}}}\\ {\bm{x}_{{\bm{q}}}}\end{matrix}\right]={\bm{x}}\,.$ (23)
In this experiment, we compare RTW with naive interpolation of the recorded data to the same length, as well as with TTW66footnotemark: 6 and NTW66footnotemark: 6 for $\mathrm{SE}(3)$ utilizing (22) and (23). Evaluation is done on two tasks, (i) following a constrained path, and (ii) repetitively drawing three circles, both shown in Figure 3.
V-E1 Following a constrained path
We instruct the robot to follow a constrained path with its end-effector, requiring precise control of position and orientation. Due to unintended time shifts within the recorded demonstration data, naive interpolation does not suffice and the resulting mean signal deviates from the desired path, resulting in a collision with the environment. TTW66footnotemark: 6, NTW66footnotemark: 6 and RTW succeed in following the constrained path without collision, however, the warping functions generated from TTW66footnotemark: 6 (15) highly compress the motion in certain sections while stretching others, resulting in a mean signal that becomes difficult to execute by the robot, as high velocities are required at the compressed sections.
V-E2 Repetitively drawing three circles
We demonstrate end-effector trajectories consisting of three circles on a whiteboard, which allows us to investigate the time warping performance for repetitive segments within provided signals. Again, naive interpolation is not sufficient. Even though both, TTW66footnotemark: 6 and NTW66footnotemark: 6, perform better, they do not align all parts of the trajectory properly. By only considering the distance between single points, TTW66footnotemark: 6 and NTW66footnotemark: 6 fail to match repetitive patterns correctly. This results in severe deviations from the desired motion at the start and end of the trajectory for both TTW66footnotemark: 6 and NTW66footnotemark: 6. Furthermore, TTW66footnotemark: 6 again suffers from fast velocity changes. Only RTW successfully aligns all parts of the demonstration data. Repetitive characteristics of the signal are correctly identified due to the Gaussian-weighted window loss (8) that considers entire segments. When executing the mean signal ${\mathbf{{\hat{y}}}}$ obtained from the proposed RTW approach, the robot’s end-effector successfully draws the three consecutive circles on the whiteboard, as shown in Figure 3.
VI Conclusion
This letter introduces Riemannian Time Warping (RTW), a novel approach that efficiently aligns multiple signals while considering the geometric structure of the Riemannian manifold the data resides in, a feature that is essential for many robotic applications. We conducted extensive experiments on real world data and with an LBR iiwa robot, proofing that RTW consistently outperforms state-of-the-art baselines on both classification and averaging tasks. Furthermore, we showed that pre-processing motion data with RTW for training a robot skill improves the learning results. The computational complexity of RTW scales linearly with the number of signals and data points, which constitutes a significant advancement over techniques based on Dynamic Programming. For many robotics problems, the optimization converges after a few seconds when utilizing a high-grade GPU.
Acknowledgments
This work was partly supported by KUKA Deutschland GmbH and the state of Bavaria through the OPERA project DIK-2107-0004/DIK0374/01.